OpenStudy (anonymous):

Order 34 x 102, 1.2 x 107, 8.11 x 10-3, and 435 from least to greatest.

OpenStudy (nottim):

so...did any work on this yet?

OpenStudy (nottim):

all that is required of you is multiplication. you're allowed using calculators right?

OpenStudy (anonymous):

34×10^2,1.2×10^7,8.11×10^−3,435 sorry lol

OpenStudy (anonymous):

3400 , 12000000, 00.811, 435 , 3400 , put them in order

OpenStudy (anonymous):

.00811, 435, 3400, 12,000,000 ?????

OpenStudy (anonymous):

ty charmingbri, i wasnt sure i understood

OpenStudy (anonymous):

435=4.35×102 34×102=3.4×103 So, now, you get 3.4×103,1.2×107,8.11×10−3,4.35×102 Rank according to the power.. smaller power of 10 -> smaller value of the number. And your welcome , medal me (:

OpenStudy (nottim):

o course ppl, try not giving the answer away. this site's for helping people, not answers

OpenStudy (anonymous):

i was typing the answers when she confirmed

OpenStudy (anonymous):

you're right , but if you look closely i actually did help him , with the problem and gave him the answer so he wouldn't fall back in the same confusion he was in but thank you anyways Tim (: aha.